# Large Language Models as Realistic Microservice Trace Generators

### v1

@@ 3.1.2 Recursive Generation
We propose to break down the task of generating a call graph into a series of recursive layer generation tasks to handle complex structures. Starting from the initial attributes, or prompt $\mathbf{c}$, the task for the model at each layer is to generate the edges originating from the Start Node specified in the prompt. The model also generates a new prompt for the next layer based on the previous layer prompt and the edges generated in the current layer. This new prompt is then re-used to condition the model’s output for the next layer. The recursive process continues until the requested attributes $\mathbf{c}$ are satisfied.
Formally, for an encoded call graph $\mathbf{t}=[\mathbf{t}_{1},\mathbf{t}_{2},\dots,\mathbf{t}_{n}]$, we partition the edges $\mathbf{t}_{i}$ into a sequence of layers $[\mathbf{t}^{1},\mathbf{t}^{2},\dots,\mathbf{t}^{l}]$, where $l\leq n$. Each layer is comprised of a sequence of edges that share the same parent (i.e., source) node, and no two edges are shared by layers. For call graph conditions $\mathbf{c}$ that describe $\mathbf{t}$, we introduce layer conditions $\mathbf{c}^{j}$, $j\in\{1,2,\dots,l+1\}$. Layer condition $\mathbf{c}^{j}$ encodes the attributes of the remaining portion of the call graph after the sequence of layers $[\mathbf{t}^{1},\mathbf{t}^{2},\dots,\mathbf{t}^{j-1}]$ has been generated, and we define $\mathbf{c}^{1}\coloneqq\mathbf{c}$ and $\mathbf{c}^{l+1}\coloneqq\emptyset$. We decompose the conditional call graph distribution as a chain of conditional layer distributions:
$p(\mathbf{t}|\mathbf{c})=\prod_{k=1}^{l}p(\mathbf{c}^{k+1},\mathbf{t}^{k}|\mathbf{c}^{k})$ (1)
In other words, the model predicts call graphs from user prompts iteratively layer-by-layer. For layer $k$ the model takes conditions $\mathbf{c}^{k}$ as input and produces the sequence of edges $\mathbf{t}^{k}$ followed by the conditions $\mathbf{c}^{k+1}$ of the next layer. The model-generated conditions $\mathbf{c}^{k+1}$ are then re-used as inputs to predict the next layer, $k+1$. Figure 2 illustrates an example of a recursively generated call graph.
@@ 3.2 Instruction Tuning
We perform supervised fine-tuning after pre-training to improve the model’s ability to generate call graphs following user instructions. Different from pre-training, we do not calculate loss for the model on the initial call graph attributes $\mathbf{c}$ (equivalent to the first layer conditions $\mathbf{c}^{1}$), which are now treated as a fixed prompt. The user can supply additional natural language instructions for the model, and in § 4.4, we provide results for two types of additional instruction. We further supplement the instructions with additional prompts, which can be programmatically generated from a template based on the user-requested attributes, to aid the model’s reasoning abilities, as detailed in § A.3. These prompts convert the attributes, including numbers, application ID strings, and other non-natural language inputs, into natural language instructions.
@@ 3.2.1 Intermediate Instructions
We find that the model often struggles to generate consistent and correct next layer conditions $\mathbf{c}^{k+1}$ based on the current layer edges $\mathbf{t}^{k}$ and conditions $\mathbf{c}^{k}$ during recursive generation. For instance, the model may generate conditions that violate physical constraints, such as assigning a higher latency to a layer than the overall call graph. Inspired by recent work demonstrating that LLM ability improves when explicitly forced to reason step-by-step (Wei et al., 2022; Nye et al., 2021), we propose including a series of natural language reasoning steps that reinforce the model’s ability to adhere to constraints. For example, we include a step-by-step calculation to (1) find the number of remaining edges based on the Num Edges attribute in $\mathbf{c}^{k}$ and the number of edges generated in $\mathbf{t}^{k}$, and (2) decide the Remaining Depth attribute in $\mathbf{c}^{k+1}$ from the same attribute in $\mathbf{c}^{k}$ (e.g., Child’s remaining depth = the current remaining depth - 1 = ...). We include these intermediate instructions immediately before the next layer conditions $\mathbf{c}^{k+1}$ during instruction fine-tuning. We give an example of these reasoning steps in § A.3.
@@ 4 Evaluation
We demonstrate the effectiveness of our method in two major aspects: (1) synthetic trace quality in terms of structural validity (§ 4.1), distribution similarity (§ 4.2), and usefulness to train and evaluate machine learning-driven microservice management tasks (§ 4.3), and (2) benefits from our use of LLMs in terms of instruction-following capabilities (§ 4.4) and downstream task performance (§ 4.5).
We initialize our model from Llama-2 7B (Touvron et al., 2023) and train with LoRA (Hu et al., 2022) on 1.36 million microservice call graph samples from the Alibaba v2022 dataset (Luo et al., 2022), corresponding to 1.1B tokens. We reserve 10% of these samples for validation. Instruction tuning datasets were created by randomly selecting 5% of the training graphs and reformatting them for instruction tuning. The training lasted four epochs, using a temperature of 0.8 and top-K of 50 for trace generation, unless otherwise specified. Further details on data preprocessing and training hyperparameters are provided in Appendix A. We compare synthetic trace quality with various structured data generation methods such as GReaT (Borisov et al., 2023) and TVAE (Xu et al., 2019), and downstream task performance with one of the state-of-the-art LLMs, Llama-3.1 405B.
4.1 Structured Reasoning with Recursive Generation and Instruction Tuning
This experiment demonstrates how recursive generation and instruction tuning with intermediate instructions enhance LLMs’ ability to accurately construct microservice call graphs. We evaluate our model by generating traces with specified num_edges and depth. A trace is deemed accurate if it correctly matches the specified num_edges and depth and adheres to all structural constraints, such as valid DAG formations and appropriate start/finish times for communications, detailed in Appendix B. We generate 50 samples for each (num_edges, depth) pair across ranges of $1\leq$ num_edges $\leq 30$ and $1\leq$ depth $\leq 6$.
@@ 4.1 Structured Reasoning with Recursive Generation and Instruction Tuning
Figure 3: Call graph generation accuracy with varying (a) edges and (b) depth in prompt using greedy sampling. The plot in (c) shows the accuracy with varying sampling temperature. Accuracy measures the fraction of generated traces that are valid and follow the initial instructions. As shown, both recursive generation and instruction tuning help to increase the validity of the synthetic traces.
Baselines.  We compare our model (recursive + instruction) to Llama-2 7B models trained on text-encoded call graphs (1) without recursive generation and tuning with intermediate instructions (baseline) and (2) with recursive generation but no instruction tuning (recursive). Both baseline models are given num_edges and depth at the start of each sample during training (see Figure 9 for an example of a training sample for the baseline model). Baselines are trained using the same hyperparameters and number of tokens as our model. For the baseline model, we represent call graph traces as the tabular data format following the method in GReaT (Borisov et al., 2023).
Results.  3(a) and 3(b) present the accuracy of generated call graphs across varying numbers of edges and depths. Generally, as complexity increases (i.e., more edges or greater depth), the baseline model’s accuracy decreases significantly—dropping below 25% for edges greater than 15 and nearing zero for depths above four. In contrast, the recursive generation model maintains higher accuracies, approximately 30% and 35%, respectively. This improved performance is attributed to the model breaking down complex generation tasks into simpler, more manageable sub-tasks.
3(c) illustrates how accuracy varies with the temperature parameter during decoding. Both models show decreased performance as the temperature increases, but the recursive model consistently outperforms the baseline, maintaining about 10% higher accuracy even at a temperature of 1. Further, instruction tuning enhances model accuracy—from 23% to 36%— by directing the model to adhere to specific generation instructions, such as the number of edges per layer, which are outlined in § A.3.
4.2 Similarity between Real and Synthetic Traces
To evaluate the quality of synthetic traces, we compare similarities between real traces from the training dataset and synthetic ones. We generate 50K call graph traces using prompts generated by the validation dataset and compare to the call graphs in the validation dataset.
@@ 4.2 Similarity between Real and Synthetic Traces
    Llama-2 7B + tabular format (GReaT (Borisov et al., 2023)): A Llama-2 7B model fine-tuned on the tabular data format of call graph traces (Same as baseline in § 4.1).
@@
    Probabilistic model: Probabilistic model based microservice call graph generators by Alibaba (Luo et al., 2021). The model is designed to follow the random distribution of different statistics, such as communication types and the number of children per depth.
@@
    TVAE (Xu et al., 2019): Tabular data generative model using VAE (Goodfellow et al., 2014). Since tabular data cannot be used to generate traces, we use the baseline only to compare distributions of popular edges. Also, to limit the training data size, we randomly choose 100k training samples from the trace dataset and use SDV (Patki et al., 2016) to train.
Distribution of Popular Calls.  Realistic synthetic traces should mirror real-world communication patterns. To assess this, we analyze the distribution of calls, defined by the attributes (Source, Destination, Communication type). 4(a) illustrates the distributions of the 100 most popular calls generated by our method and the baselines, limited to the top 30 due to space constraints.
@@
Heavy-hitter Prediction.  The capability to generate heavy-hitter microservices—defined as top-$K$ microservices triggered in a sequence of call graphs—is critical for tasks such as resource optimization and anomaly detection in microservice management. In this experiment, we select 1K traces from the validation dataset and create instructions consisting of a service ID and call graph attributes such as depth and the number of edges. These instructions guide the synthetic trace generation for both the baseline and our models. We evaluate the accuracy by comparing the top-$K$ microservices between the synthetic and validation traces over 20 runs.
@@
4(b) illustrates the accuracy for varying $K$ values, ranging from 10 to 500. Our method demonstrates robust performance, maintaining over 90% accuracy for $K\leq 50$ and 65% at $K$=500. The model trained with the GReaT method also shows robust performance, but slightly worse performance with larger $K$ values. We believe that the performance gap results from the lack of capability to generate complex structures (§ 4.1), which can affect the distribution of generated traces. On the other hand, the probabilistic model starts at 59% accuracy for $K$=10 and declines to 23% at $K$=500, showcasing our method’s capability to capture and replicate heavy-hitter dynamics in synthetic traces.
Additional evaluation results on the similarity of microservice branching (in-degree and out-degree) and response time distributions can be found in § D.3.
4.3 Usefulness of Synthetic Data as ML Training data
@@
The synthetic dataset is intended to serve as a substitute for real data in the training process. Thus, we assess how well state-of-the-art microservice management tasks for critical component extraction in FIRM (Qiu et al., 2020) and anomaly detection in TraceVAE (Xie et al., 2023), which use machine learning (ML) models, perform when the models are trained on the synthetic datasets. Specifically, the ML models are evaluated using real test data, and their results are compared to their original performance when trained on the real training dataset.
@@
When choosing training data, we first select a subset of traces from real data and label them with corresponding conditions (e.g., critical microservices). To have similar distributions between synthetic and real traces, we extract instructions from the selected real traces and use them to generate synthetic traces. We do not include invalid call graphs using the same accuracy metrics in § 4.1. We train each downstream task using 5K call graphs and evaluate it using 2K call graphs for testing. We consistently use the same test dataset derived from real data. We use the default hyperparameters set by FIRM and TraceVAE (Qiu et al., 2020; Xie et al., 2023), and set labels based on 99-percentile latencies. We run each experiment 5 times, varying the random seeds or test datasets, and report the average performace for the following two evaluation tasks. We use synthetic traces generated by GReaT and the Alibaba probabilistic model as baselines for comparing the performance of ML models.
@@
Critical Component Extraction.  For efficient resource management, FIRM (Qiu et al., 2020) predicts critical components (microservices likely to violate service level objectives (SLO)) from call graphs with support vector machines (SVMs) using latency-related features and determines additional resource types and amounts for the critical components. For our evaluation, we train SVMs to detect critical components using two popular applications (apps A and B) from our trace dataset. For each application, we randomly sample call graphs and train two SVMs: one with real data and one with synthetic data generated by our fine-tuned model. Figure 5 shows the accuracy of the models when evaluated on real call graphs from the test set. SVMs trained on synthetic data perform similarly to those trained on real data, differing by less than 1.5 percentage point. SVMs trained on synthetic traces from baselines exhibit an accuracy gap ranging from 6 to 81 percent points.
@@
Anomaly Detection.  For operators to efficiently diagnose system failures, anomaly detection models predict whether microservice call graphs include anomalous characteristics like irregular graph structure or time. We assess our synthetic data quality using TraceVAE (Xie et al., 2023), a variational autoencoder (VAE) model that detects anomalous microservices in terms of time consumption. We train TraceVAE models using real and synthetic trace data, similar to our previous experiment. Figure 5 reports ROC AUC for the models evaluated on real test data. Again, we see that training on synthetic data generated by our method consistently yields results comparable to those obtained from real data.
In conclusion, the synthetic traces from our method demonstrate similar performance to real traces and have the potential to be leveraged in various use cases, as demonstrated by the above evaluation tasks. We attribute our method’s comparable performance with real data to its ability to generate complex structures and capture diverse characteristics. We also find similar results with two additional classification tasks by fine-tuning Llama-2 7B models. We describe the tasks and results in § D.4.
@@ 4.4 Instruction-following Capability
Enabling users to specify desired characteristics of synthetic data is crucial for trace generators. We assess our instruction-tuned model’s capacity to reflect user-requested attributes in the generated traces accurately. We evaluate the model’s ability to produce call graphs featuring specific attributes (high latency and uncommon communications). Additionally, we explore the model’s performance when prompted with a combination of these attributes not present in the training data.
@@
    High Latency: Instructions specify that call graphs should exhibit latencies above the 90th percentile (p90) of the training dataset’s latency distribution, which varies by service. For example, the instruction might read: Build a call graph with high latency.
@@
    Uncommon Communications: Instructions indicate that the call graph layer should include a communication occurring in less than 10% of the training data. An instruction example is: Include an edge from (SRC) to (DEST) with (TYPE) communication type.
We intentionally avoid combining these specific instructions in training samples to test the model’s response to novel instruction combinations during inference.
Results.  Figure 7 presents the instruction-following accuracy for high latency and uncommon communication. We assessed this by filtering 1K validation instructions to see how many generated call graphs met the defined criteria (e.g., exceeding p90 latency). We also compared these results against outputs generated without specific instructions to evaluate the impact of tailored prompts on model performance.
Additionally, we examine the model’s performance when both instructions are combined in prompts, a scenario not covered in the training data. The model’s ability to satisfy both conditions simultaneously, despite not being explicitly trained to do so, is detailed in the right of Figure 7. Higher accuracy in scenarios without specific instructions often results from inherent biases in attributes like service ID or the number of edges, which may align with the desired user outcomes.
Figure 6: Instruction-following accuracy (%).
@@
4.5 Adapting Models to Solve Downstream Tasks
We extend our evaluation beyond generating synthetic traces, demonstrating the utility of our pre-trained model in performing downstream tasks related to microservice traces. The trace pre-trained model is adapted to each downstream task through additional fine-tuning. We focus on scenarios where partial information from distributed environment traces is available, emphasizing the challenges posed by incomplete data. This section compares our fine-tuned model with the standard Llama-2 7B, which lacks specific training on call graph data, and with Llama-3.1 405B by providing task descriptions and up to 16 examples in prompts (i.e., in-context learning (Brown et al., 2020)), to highlight the importance of domain-specific training.
Predicting Uncommon Communication Patterns.  The task is to predict uncommon communication patterns (as in § 4.4) based on the first 10 lines of a trace. We train the original Llama model and our adapted model for this binary classification task on 15K samples. Each sample’s prompt comprised the first 10 edges of a real trace, with binary labels indicating the presence of uncommon communication patterns in the subsequent trace sections.
Results, detailed in the left of Figure 7, indicate that the original Llama-2 7B model achieves only 60.6% accuracy, suggesting insufficient training for recognizing uncommon patterns. Additionally, in-context learning with Llama-3.1 405B shows lower accuracy (45.6%), indicating larger models trained on typical internet data lack capability to solve tasks from domain-specific data. In contrast, our model achieves 76.8% accuracy, demonstrating its enhanced capability to interpret and predict based on partial trace data.
Infilling Missing Data.  Missing data is common in large-scale trace logging, such as in Alibaba’s microservice call graphs, where 67% of traces contain missing values (Huye et al., 2024). This task focuses on fine-tuning our model to accurately infill missing data in microservice call graphs, considering partial information. Specifically, we conduct two separate experiments on infilling (1) a missing attribute and (2) a missing call connecting two layers.

### v2

@@ 3.1.2 Recursive Generation
We propose to break down the task of generating a call graph into a series of recursive layer generation tasks to handle complex structures. Starting from the initial attributes, or prompt $\mathbf{c}$, the task for the model at each layer is to generate the edges originating from the Start Node specified in the prompt. The model also generates a new prompt for the next layer based on the previous layer prompt and the edges generated in the current layer. This new prompt is then re-used to condition the model’s output for the next layer. The recursive process continues until the requested attributes $\mathbf{c}$ are met.
Formally, for an encoded call graph $\mathbf{t}=[\mathbf{t}_{1},\mathbf{t}_{2},\dots,\mathbf{t}_{n}]$, we partition the edges $\mathbf{t}_{i}$ into a sequence of layers $[\mathbf{t}^{1},\mathbf{t}^{2},\dots,\mathbf{t}^{l}]$, where $l\leq n$. Each layer consists of a sequence of edges that share the same parent (i.e., source) node, ensuring that no edge appears in multiple layers. For call graph conditions $\mathbf{c}$ that describe $\mathbf{t}$, we introduce layer conditions $\mathbf{c}^{j}$, $j\in\{1,2,\dots,l+1\}$. Layer condition $\mathbf{c}^{j}$ encodes the attributes of the remaining portion of the call graph after the sequence of layers $[\mathbf{t}^{1},\mathbf{t}^{2},\dots,\mathbf{t}^{j-1}]$ has been generated, and we define $\mathbf{c}^{1}\coloneqq\mathbf{c}$ and $\mathbf{c}^{l+1}\coloneqq\emptyset$. We decompose the conditional call graph distribution as a chain of conditional layer distributions: $p(\mathbf{t}|\mathbf{c})=\prod_{k=1}^{l}p(\mathbf{c}^{k+1},\mathbf{t}^{k}|\mathbf{c}^{k})$. In other words, the model predicts call graphs from user prompts iteratively layer-by-layer. For layer $k$ the model takes $\mathbf{c}^{k}$ as input and produces the sequence of edges $\mathbf{t}^{k}$ followed by the conditions $\mathbf{c}^{k+1}$ of the next layer. The process continues recursively, using $\mathbf{c}^{k+1}$ to predict the next layer, $k+1$. Figure 2 illustrates an example of this recursive generation.
@@ 3.2 Instruction Tuning
After pre-training, we perform supervised fine-tuning to enhance the model’s ability to generate call graphs based on user instructions. Unlike pre-training, we exclude the initial call graph attributes $\mathbf{c}$ (equivalent to the first-layer conditions $\mathbf{c}^{1}$) from the loss computation, treating them as a fixed prompt. Users can provide additional instructions, and § 4.4 presents results for two instruction types. To further aid reasoning, we supplement instructions with programmatically generated prompts that convert numerical and non-linguistic attributes (e.g., application IDs) into natural language, as detailed in § B.3.
@@ 3.2.1 Intermediate Instructions
The model often struggles to generate consistent next-layer conditions $\mathbf{c}^{k+1}$ based on the current layer’s edges $\mathbf{t}^{k}$ and conditions $\mathbf{c}^{k}$, sometimes violating physical constraints (e.g., assigning a layer higher latency than the overall call graph). Inspired by work showing that LLMs improve with explicit step-by-step reasoning (Wei et al., 2022; Nye et al., 2021), we introduce natural language reasoning steps to reinforce constraint adherence. For example, we (1) compute remaining edges from the Num Edges attribute in $\mathbf{c}^{k}$ and edges in $\mathbf{t}^{k}$, and (2) derive the Remaining Depth in $\mathbf{c}^{k+1}$ as Child’s remaining depth = current remaining depth - 1 = .... These intermediate instructions are inserted before $\mathbf{c}^{k+1}$ during instruction fine-tuning. We give an example of these steps in § B.3.
@@ 4 Evaluation
We thoroughly demonstrate the effectiveness of our method in two major aspects: (1) synthetic trace quality in terms of structural validity (§ 4.1), distribution similarity (§ 4.2), and usefulness to train and evaluate machine learning-driven microservice management tasks (§ 4.3), and (2) benefits from our use of LLMs in terms of instruction-following capabilities (§ 4.4) and trace-related downstream task performance (§ 4.5).
We initialize our model from Llama-2 7B (Touvron et al., 2023) and train with LoRA (Hu et al., 2022) on 1.36 million microservice call graph samples from the Alibaba v2022 dataset (Luo et al., 2022), corresponding to 1.1B tokens. Further details on data preprocessing and training hyperparameters are provided in Appendix B. We compare synthetic trace quality with various structured data generation methods such as GReaT (Borisov et al., 2023) and TVAE (Xu et al., 2019), and downstream task performance with one of the state-of-the-art LLMs, Llama-3.1 405B (Dubey et al., 2024).
@@
Figure 3: Call graph generation accuracy with varying (a) edges and (b) depth in prompt using greedy sampling. (c) shows the accuracy with varying sampling temperature. Accuracy measures the fraction of generated traces that are valid and follow the initial instructions. As shown, both recursive generation and instruction tuning improve the accuracy of the synthetic traces.
4.1 Structured Reasoning Accuracy
This experiment demonstrates how recursive generation and instruction tuning with intermediate instructions enhance LLMs’ ability to accurately construct microservice call graphs. We evaluate our model by generating traces with specified num_edges and depth. A trace is deemed accurate if it correctly matches the specified num_edges and depth and adheres to all structural constraints, such as valid DAG formations and appropriate start/finish times for communications, detailed in Appendix C. We generate 50 samples for each (num_edges, depth) pair across ranges of $1\leq$ num_edges $\leq 30$ and $1\leq$ depth $\leq 6$.
Baselines.  We compare our model (recursive + instruction) to Llama-2 7B models trained on text-encoded call graphs (1) without recursive generation and tuning with intermediate instructions (baseline) and (2) with recursive generation but no instruction tuning (recursive). Both baseline models are given num_edges and depth as inputs during training (see Figure 9 for a training data example of the baseline model). Baselines are trained using the same hyperparameters and number of tokens as our model. The baseline model follows GReaT (Borisov et al., 2023), representing call graph traces as the tabular data format.
Results.  3(a) and 3(b) present the accuracy of generated call graphs across varying numbers of edges and depths. Generally, as complexity increases (i.e., more edges or greater depth), the baseline model’s accuracy decreases significantly—dropping below 25% for edges $>$ 15 and nearing zero for depths $>$ 4. In contrast, the recursive generation model maintains higher accuracies, by approximately 30% and 35%, respectively. This improved performance is attributed to the model breaking down complex generation tasks into simpler, more manageable sub-tasks.
3(c) illustrates how decoding temperature affects accuracy. Both models show decreased performance as the temperature increases, but the recursive model consistently outperforms the baseline, maintaining about 10% higher accuracy even at a temperature of 1. Further, instruction tuning enhances model accuracy by 23% to 36% by directing the model to adhere to specific generation instructions, such as the number of edges per layer, which are outlined in § B.3.
More results on accuracy with varying model sizes and memorization are in § D.2 and § D.3.
4.2 Similarity of Real and Synthetic Traces
To assess the quality of synthetic traces, we compare them to real traces from the validation dataset. We generate 50K call graphs using prompts derived from the validation set and measure their similarity to the original traces.
@@ 4.2 Similarity of Real and Synthetic Traces
    GReaT (Borisov et al., 2023) (Llama-2 7B + tabular format): A Llama-2 7B model fine-tuned on the tabular data format of call graph traces (Same as baseline in § 4.1).
@@
    Probabilistic model: A call graph generator from Alibaba (Luo et al., 2021) that samples graph structures based on statistical distributions, such as communication types and the number of child nodes per depth.
@@
    TVAE (Xu et al., 2019): A VAE-based tabular data generator (Goodfellow et al., 2014). Since it cannot directly generate traces, we use it to compare edge distributions. Training is limited to 100K randomly selected samples.
Distribution of Popular Calls.  Realistic synthetic traces should mirror real-world communication patterns. To assess this, we analyze the distribution of calls, defined by (Source, Destination, Communication type). 4(a) compares the 100 most popular calls generated by our method and the baselines, displaying the top 30 for clarity.
@@
Heavy-hitter Prediction.  Generating top-$K$ heavy-hitter microservices—those most frequently triggered in a sequence of call graphs—is crucial for resource optimization and anomaly detection in microservice management. In this experiment, we select 1K validation traces and create instructions with a service ID and call graph attributes like depth and #edges to guide trace generation for both our model and the baseline. Similarity is evaluated by comparing the top-$K$ microservices between synthetic and validation traces over 20 runs.
@@
4(b) illustrates the similarity for varying $K$ values, from 10 to 500. Our method achieves over 90% accuracy for $K\leq 50$ and 65% at $K$=500, demonstrating robust performance. The model trained with the GReaT method also shows robust performance, but slightly worse performance with larger $K$ values. We believe that the performance gap results from the lack of capability to generate complex structures (§ 4.1), which impacts trace distribution. In contrast, the probabilistic model starts at 59% accuracy for $K$=10 and declines to 23% at $K$=500, showcasing our method’s capability to capture and replicate heavy-hitter dynamics.
Additional evaluation on branching (in/out-degree) and latency distributions is in § D.4.
4.3 Synthetic Data as Training data for ML-Driven Microservice Management
@@
Synthetic datasets can be used as a substitute for scarce real data in the training process for ML-driven microservice management tasks. Thus, we assess how well microservice management tasks using ML models for critical component extraction in FIRM (Qiu et al., 2020) and anomaly detection in TraceVAE (Xie et al., 2023) perform when the models are trained on the synthetic datasets. The ML models are evaluated using real test data, and their results are compared to their original performance when trained on the real training dataset.
@@
When choosing training data, we select a subset of traces from real data and label them with corresponding conditions (e.g., critical microservices). Then, we extract instructions from real data and use them to generate synthetic traces. We exclude invalid call graphs using the same accuracy metrics in § 4.1 before training. We train the models on 5K synthetic call graphs and test on 2K real call graphs, using the same test dataset across all experiments. For baselines, we use synthetic traces generated by GReaT and the Alibaba probabilistic model. Each experiment is run 5 times, varying random seeds, and results are averaged.
@@
Critical Component Extraction.  FIRM (Qiu et al., 2020) predicts critical components (microservices likely to violate service level objectives (SLO)) using support vector machines (SVMs) trained on latency-related features from call graphs. For our evaluation, we train SVMs to detect critical components using two popular applications (apps A and B) from our trace dataset. For each application, we randomly sample call graphs and train two SVMs: one with real data and one with synthetic data generated by our fine-tuned model. As shown in Figure 5, SVMs trained on synthetic data achieve near-identical accuracy to those trained on real data, with a difference of less than 1.5 percentage points. In contrast, SVMs trained on synthetic traces from baselines show a performance gap of 6 to 81 percentage points.
@@
Anomaly Detection.  For operators to efficiently diagnose system failures, anomaly detection models predict whether microservice call graphs include anomalous characteristics like irregular graph structure or time. We assess our synthetic data quality using TraceVAE (Xie et al., 2023), a variational autoencoder (VAE) model that detects anomalous microservices in terms of time consumption. We train TraceVAE models using real and synthetic trace data, similar to our previous experiment. Figure 5 shows that models trained on synthetic traces from our method yield results comparable to those trained on real data, as measured by ROC AUC.
We obtained similar results with two other classification tasks using fine-tuned Llama-2 7B models, as detailed in § D.5.
@@ 4.4 Instruction-following Capability
Enabling users to specify desired characteristics of synthetic data is crucial for trace generation. Such "custom" traces are useful to study corner cases and debug microservice management techniques. We assess our instruction-tuned model’s capacity to accurately generate call graphs with specified attributes, such as high latency and rare communication types. We also explore the model’s performance when prompted with combinations of these attributes that were not included in the training data.
@@
    High Latency: Instructions specify that call graphs should exhibit latencies above the 90th percentile (p90) of the training dataset’s distribution, varying by service. For example: Build a call graph with high latency.
@@
    Uncommon Communications: Instructions indicate that the call graph layer should include a communication occurring in less than 10% of the training data. For example: Include an edge from (SRC) to (DEST) with (TYPE) communication type.
We avoid combining these specific instructions in training samples to test the model’s response to novel instruction combinations during inference.
Results.  Figure 7 presents the instruction-following accuracy for high latency and uncommon communication. We assessed this by filtering 1K validation instructions to see how many generated call graphs met the defined criteria (e.g., exceeding p90 latency). We also compared these results against outputs generated without specific instructions to assess the impact of tailored prompts.
Additionally, we evaluate the model’s performance when both instructions were combined, a scenario not covered during training. The model’s ability to satisfy both conditions simultaneously, despite not being explicitly trained to do so, is detailed in the right of Figure 7. Higher accuracy in the absence of instructions may arise from inherent biases, such as those related to service ID or edge count, which align with the desired outcomes.
Figure 6: Instruction following accuracy (%).
@@
4.5 Adapting Models for Trace-Related Downstream Tasks
We extend our evaluation beyond generating synthetic traces, demonstrating the utility of our pre-trained model in performing downstream tasks related to microservice traces. The trace pre-trained model is adapted to each downstream task through additional fine-tuning. We focus on scenarios where partial information from distributed environment traces is available, emphasizing the challenges posed by incomplete data. This section compares our fine-tuned model with the standard Llama-2 7B, which lacks specific training on call graph data, and with Llama-3.1 405B by providing task descriptions and up to 16 examples in prompts (i.e., in-context learning (Brown et al., 2020)), to highlight the need for domain-specific training.
Predicting Uncommon Communications.  The task is to predict uncommon communication patterns (as in § 4.4) based on the first 10 lines of a trace. For fine-tuning, we adapt both the original Llama-2 7B and our trace pretrained model to this binary classification task using 15K samples. Each sample’s prompt comprised the first 10 edges of a real trace, with binary labels indicating the presence of uncommon communication patterns in the subsequent trace sections.
As shown on the left side of Figure 7, the original Llama-2 model achieves only 60.6% accuracy, indicating insufficient training for recognizing uncommon patterns. Additionally, in-context learning with Llama-3.1 405B shows lower accuracy (45.6%), suggesting that larger models trained on general internet data struggle with domain-specific tasks. In contrast, our model achieves 76.8% accuracy, demonstrating its enhanced capability to interpret and predict based on partial trace data.
Infilling Missing Data.  Missing data is common in large-scale trace logging, such as in Alibaba’s microservice call graphs, where 67% of traces contain missing values (Huye et al., 2024). This task focuses on fine-tuning our model to accurately infill missing data in microservice call graphs, considering partial information. Specifically, we conduct two experiments on infilling (1) a missing attribute and (2) a missing call connecting two layers.